OpenStudy (anonymous):

Solve by the elimination method: x + y = 14 4x - y = 31

jhonyy9 (jhonyy9):

assume it x+y=14 4x-y=31 ----------- 5x =45 x = 45/5 = 9 x=9 than y=14-9=5 x=9 y=5